In each of 4 races, the Democrats have a 60% chance of winning. Assuming that the races are independent of each other., what is the probability that:




1
Expert's answer
2021-10-06T12:32:01-0400

Solution:

We need to find the probability that the democrats will win at least 1 race.

Let the number of wins out of 4 races be modelled as "X".

X"\sim Bin(n,p)"

n=4, p=0.60

"P(X\ge 1)=1-P(X=0)
\\=1-(0.6)^0(0.4)^4
\\=0.9744"